OpenStudy (anonymous):

Assume that a>0. Differentiate the function at a^t

OpenStudy (samiam):

ta^(t-1)???

OpenStudy (anonymous):

well I thought that too , but no because a is always used to identify an actual number and not a variable

OpenStudy (samiam):

oh whoops didnt notice that give me a sec so the variable is the t???

OpenStudy (samiam):

so assuming that a is a value and t is the variable: ln(a)a^t

OpenStudy (samiam):

this is one of the rules

OpenStudy (anonymous):

thats what I was thinking also let me check thanks

OpenStudy (samiam):

hey r u there?

OpenStudy (samiam):

sorry i just lost connection for a sec

OpenStudy (samiam):

so was it correct?

OpenStudy (anonymous):

Yes, sorry lost connection as well. I thought it was something to that effect, thanks!

OpenStudy (samiam):

hey u r still here. thought u were long gone :D

OpenStudy (samiam):

ok bye for now gonna go study calc

OpenStudy (jamesj):

Yes, and you can prove it this way: \[ a^x = (e^{\ln a})^x \] because a = e^(ln a) Now, \[ (e^{\ln a})^x = e^{(\ln a)x} \] therefore \[ \frac{d \ }{dx} a^x = \frac{d \ }{dx} e^{(\ln a)x} = \ln a .e^{(\ln a)x} \] because ln a is just a constant. Finally, \[ \ln a .e^{(\ln a)x} = \ln a . a^x \] therefore \[ \frac{d \ }{dx} a^x = \ln a . a^x \]

OpenStudy (samiam):

james always gives the best answer. TRUST JAMES :D